Calculate the bond price for maturity 4 years with face (nominal) value 100, annual coupons 5, and continous compounding with interest a) 8% and b) 5%. What do you notice?
Answer #1

For continuous compounding, A= P*e^(r*t)

So, P= A/e^(r*t)

a)Given, r= 8%

P= 100/e^(0.08*4)=100/1.38= 72.46

b) r=5%

P= 100/e^(0.05*4)= 100/1.22= 81.96

So, as the interest rate decreases the present value of the bond is higher.